# Compressible and immiscible fluids with arbitrary density ratios: diffuse interface theory

### v2

@@ VI.2 Kinetic equations
The coefficient $\varsigma_{i}$ is the mobility.
@@
The conservative force at the right hand side is responsible for the exchange of the potential energy $\mathcal{F}+\mathcal{P}$ with the kinetic energy $\mathcal{K}$. This result is nothing but Newton’s law of momentum balance. Here, a more clear physical meaning of the force is presented; the force term consists of the pressure $p$ and the stress tensor $\underline{\underline{\bm{\Theta}}}$. The stress tensor is due to the gradient term in the potential energy $\mathcal{F}$, while the pressure is contributed by the associated free energy. Note that a viscous dissipation term cannot be directly added to the momentum conservation equation, as there is no quantity corresponding to “viscous energy”.
The only term left in the time evolution of system energy, III, is $\int_{\Omega}\bm{\zeta}\cdot\dot{\mathbf{u}}\mathrm{d}\Omega$. By applying the same dissipation principle as done for the scalar field in Eq. (65), we obtain the dissipation equation for the velocity as
@@
Unlike the dissipation rule for the scalar variable $\phi_{i}$ with the associated potential $\nabla\mu_{i}$ (see Eq. (65)), three terms $\nabla\bm{\zeta}$, $\nabla\bm{\zeta}^{T}$, and $(\nabla\cdot\bm{\zeta})\mathbf{I}$ have to be taken into account for the dissipation of a vector $\mathbf{u}$.
@@
$\nabla\cdot\tau_{1}\nabla\bm{\zeta}=\nabla\cdot\tau_{1}\nabla(\partial\mathcal{K}/\partial\mathbf{u});\ \bm{\zeta}=\partial\mathcal{K}/\partial\mathbf{u};$ (69)
$\nabla\cdot\varsigma\nabla\mu=\nabla\cdot\varsigma\nabla(\partial\mathcal{F}/\partial\mathbf{\phi});\ \mu=\partial\mathcal{F}/\partial\phi.$ (70)
@@
where $\tau_{0}$ is a positive mobility. The non-conserved form is known as the Allen-Cahn type dissipation following the gradient descent path, being consistent with the Langevin equation.
VI.3 The problem of the conventional kinetic energy formulation
Conventional formulation for the kinetic energy is $\tilde{\mathcal{K}}=\int_{\Omega}\frac{1}{2}\rho\mathbf{u}^{2}\mathrm{d}\Omega$. In this case, the time derivative of the kinetic energy reads
$\frac{\mathrm{d}{\tilde{\mathcal{K}}}}{\mathrm{d}t}=\int_{\Omega}\bigg(\frac{1}{2}\frac{\mathrm{d}\rho}{\mathrm{d}t}\mathbf{u}^{2}+\rho\mathbf{u}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}\bigg)\mathrm{d}\Omega.$ (72)
A noteworthy shortcoming of Eq. (72) is the missing information of the momentum balance as well as the transformation between the potential and kinetic energies. Moreover, the first part in Eq. (72) requires a prior knowledge about the time evolution of the density, which is unknown for two immiscible phases. The density evolution can be avoided if the kinetic energy is written as $\rho\mathbf{u}\cdot\mathbf{u}=\bm{\zeta}\cdot\mathbf{u}$; in this case, the time evolution of the momentum $\bm{\zeta}$ as a whole term is sufficient, subjected to the momentum conservation. Noteworthily, the so-called continuity equation $\partial_{t}\rho+\nabla\cdot(\rho\mathbf{u})=0$ cannot be applied here for density evolution. The reason is as follows: (I) The classic continuity equation is derived for pure fluid, where there is no density variation in space. (II) When considering two immiscible fluids, the classic continuity equation cannot be straightforwardly adopted, since the density varies across the interface. In this case, the mass conservation and energy dissipation are not independent. Suffice it to say, one cannot directly apply a mass conservation without considering the energy dissipation across the interface. (III) Within the interface of two immiscible fluids, one cannot use the condition $\nabla\cdot\mathbf{u}=0$ which is used for pure fluids. The divergence free condition $\nabla\cdot\mathbf{u}=0$ implies a uniform excess volume everywhere in the whole domain $\Omega$ and that the mass is uniformly distributed in the speed of sound. In other words, when the fluid velocity is much less than the speed of sound, the density almost has no variation in space; only in this case, the condition $\nabla\cdot\mathbf{u}=0$ can be applied. For pure homogeneous fluids, the condition $\nabla\cdot\mathbf{u}=0$ is used to solve the unknown pressure $p_{0}$, which is homogeneous and can be considered as a reference value (Fig. 3). However, for two immiscible fluids, the pressure is inhomogeneous across the interface, which cannot be solved simply by $\nabla\cdot\mathbf{u}=0$ due to the density variation.
VI.4 Density evolution equation and the generalization of Bernoulli’s law
@@
As demonstrated in the following, Eq. (73) provides a generalization of Bernoulli’s law.
@@
$\mathbf{u}^{2}\nabla\rho=-[\nabla\cdot(p\mathbf{I}+\underline{\underline{\bm{\Theta}}})].$ (74)
By considering a one-dimensional setup and using the relation $\mathrm{d}(\rho u^{2})=u^{2}\mathrm{d}\rho+2\rho udu$, Eq. (74) is further simplified as
$\mathrm{d}(p+\rho u^{2})=2u\mathrm{d}\rho.$ (75)
This result is nothing but a generalized Bernoulli’s equation. When there is no density variation, namely, $\mathrm{d}(p+\rho u^{2})=0$, we have
@@
which is Bernoulli’s principle. However, when density varies, $\mathrm{d}\rho\neq 0$, the associated energy is changed as well; in this case, the classic Bernoulli’s equation loses its validity.
@@ VI.5 Replication of the Navier-Stokes equations
$\frac{1}{2}\rho\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=\rho\{\nabla\cdot\tau_{1}[\nabla(\rho\mathbf{u})+\nabla(\rho\mathbf{u})^{T}]+\nabla\cdot\tau_{2}\nabla\cdot(\rho\mathbf{u})\mathbf{I}\}=\nabla\cdot\eta(\nabla\mathbf{u}+\nabla\mathbf{u}^{T})+\nabla\cdot\nu\nabla\cdot\mathbf{u})\mathbf{I}.$ (79)
@@
We stress that viscous dissipation inherently contradicts the incompressibility condition. When viscous dissipation occurs, density cannot remain constant, as dissipative processes—such as heat release, light emission, or electrical discharge—introduce entropy changes that modify the density. However, under dissipation-free conditions (e.g., isentropic flow), density can reasonably be treated as constant. In this way, the Navier-Stokes equation has the conflicts between viscous dissipation and invariance of density (incompressibility). In fact, Eq. (78) cannot be summed with Eq. (79) because:
@@
The dissipation formulation, Eq. (79) is particularly chosen to be consistent with the Newton viscous stress tensor to dissipate the velocity and the associated energy as
$\int_{\Omega}\rho\mathbf{u}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}\mathrm{d}\Omega=-\tau_{1}\int_{\Omega}\nabla(\rho\mathbf{u}):\nabla(\rho\mathbf{u})\mathrm{d}\Omega-\tau_{2}\int_{\Omega}[\nabla\cdot(\rho\mathbf{u})]^{2}\mathrm{d}\Omega\leq 0.$ (81)
Other dissipation formations for $\mathrm{d}\mathbf{u}/\mathrm{d}t$ are also possible as long as the integration
$\int_{\Omega}\rho\mathbf{u}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}\mathrm{d}\Omega$ (82)
is not greater than 0.
@@ VI.6 Remarks
The key point of the present theory is the energy minimization based on an alternative formulation for the kinetic energy, $\rho\mathbf{u}\cdot\mathbf{u}$, rather than the conventional formulation, $\frac{1}{2}\rho\mathbf{u}^{2}$. In our derivation, the dissipation of the kinetic energy is divided into two parts. The first part is the momentum conservation, $\mathrm{d}(\rho\mathbf{u})/\mathrm{d}t$ which changes not only the fluid velocity $\mathbf{u}$ but also the density $\rho$. This derivation differs from Euler and Navier-Stokes equations where the momentum variation is only caused by the evolution of the fluid velocity. The second part of the kinetic energy dissipation is the velocity dissipation due to the difference in the density, viscosity, and velocity, as explained in Fig. 4III.
Another vital point of the current work is the pressure energy, the formulation of which is consistent with many pioneers Chen and Doolen (1998). The physical meaning of the pressure has been elucidated in Fig. 3. The minimization of the pressure energy leads to the so-called grand potential model Wang et al. (2012). In the lattice model, the free energy and the pressure can be explicitly formulated according to the theory of thermodynamics, such as Landau potential. However, for non-lattice models, the pressure cannot be explicitly identified. This case is the consideration in Euler and Navier-Stokes equations, where the unknown pressure is solved by the divergence free of the velocity for incompressible flow. For liquid and solid phases, the lattice model can be applied because of the regularized order of the structure within the short range. For gas phases, the lattice model loses its validity due to the disorder of the structure; one has to use the equation of state to formulate the partial pressure of the gas phase Wöhrwag et al. (2018); Ledesma-Aguilar et al. (2014).
In the current work, we consider the linear momentum $\bm{\zeta}=\rho\mathbf{u}$. Non-linear terms related to fluid velocity may be accounted for the dissipation if we consider non-linear momentum, for example, $\bm{\zeta}\propto\mathbf{u}^{n}$ or $\bm{\zeta}\propto(\nabla\mathbf{u})^{n}$. The non-linear momentum not only modifies the velocity dissipation but also changes the momentum balance, which may be applied for understanding highly nonlinear effects, such as turbulence. These specialized cases are out of the scope of the present discussion.
@@
The evolution of the velocity corresponds to the dissipation according to Eq. (67). The local velocity can increase (acceleration) or decrease (deceleration), but overall the kinetic energy $\mathcal{K}$ decreases with time in a closed system. Once dissipation takes place, i.e., $\mathrm{d}\mathbf{u}/\mathrm{d}t\neq 0$, the density at the interface cannot be assumed to be a constant; in this case, the density evolves with time. The density evolution is especially important when the fluid velocity is close to the speed of sound, i.e., high Mach number fluid. Note that in reality, dissipation cannot be avoided, since entropy always increases with time.
@@
The first part corresponds to the momentum balance. The second part is associated with the velocity dissipation. Overall, $\mathrm{d}\mathcal{K}/\mathrm{d}t\leq 0$, which is the key basis of the present finding.
@@ VII.1 Sound speed
which is consistent with the results of the perturbation analysis Landau and Lifshitz (2013). Note that two conditions have been applied to derive the sound speed: $\delta q\approx 0$ (see appendix) and $\eta\approx 0$, known as isentropic condition (adiabatic and reversible). For one component fluid, the Landau potential is expressed as $f-\sum_{i}\mu_{i}\phi_{i}=f$, where the free energy $f$ is written as $f=\theta-Ts$. In the case of isentropic condition, the free energy is mainly contributed by the internal energy, which consists of microscopic kinetic energy $\varrho$ and microscopic potential energy $\xi$,
@@ VII.2 Equation of the state (EOS)
which is the EOS of ideal gas with a factor, $\ln\rho$. Note that the thermodynamic pressure $\Psi$ should be used instead of $p$. The term $\ln\rho$ is analogous to the entropy contribution, and this result is consistent with previous work on the free energy formulation of a water-gas system Ledesma-Aguilar et al. (2014). By dividing both sides of Eq. (96) by the molar volume, we obtain $\Psi\sim\rho\ln\rho$, as used in Ref. Ledesma-Aguilar et al. (2014).
@@ VII.3 Water-air interface
$u_{0}^{2}\frac{\mathrm{d}\rho}{\mathrm{d}x}=-\frac{\mathrm{d}p}{\mathrm{d}x},$ (105)
@@ Acknowledgements.
This research is supported by VirtMat project of the Helmholtz association, as part of the programme “MSE-materials science and engineering” number-43.31.01. F.W. proposed the concept, carried out the theoretical derivations, and prepared the manuscript. B.N. contributed to securing the funding.
Appendix A Gibbs-Duhem equation and the kinetic energy
In the case of incompressible fluids, a single variable—either $\phi$ or $\rho$—is sufficient to characterize the state of the system. For compressible fluids with nearly identical densities, a single variable likewise suffices to adequately describe the system, as demonstrated in Section III. Here, we illustrate the Gibbs–Duhem equation by taking $\phi$ as the state variable. According to the first law of thermodynamics, the change in the internal energy density $\theta$ of a unit volume element $v$ can be expressed as
$\mathrm{d}\theta=\delta q-p\mathrm{d}v+\sum_{i}\mu_{i}\mathrm{d}\phi_{i}.$ (112)
The first term corresponds to heat transfer, while the second and third terms represent mechanical and chemical work, respectively. Typically, the timescale associated with heat transfer, represented by $\delta q$, is much longer than that of mass transport, as thermal diffusivity exceeds mass diffusivity by at least three orders of magnitude. Consequently, we assume an isothermal condition on the timescale of mass transport and adopt $\delta q\approx 0$ in the current work. Nonetheless, the temperature evolution equation is derived in Section V to provide a complete description of the system.
Combining Eq. (112) with the total derivative $\mathrm{d}\theta=-p\mathrm{d}v-v\mathrm{d}p+\sum_{i}\mu_{i}\mathrm{d}\phi_{i}+\sum_{i}\phi_{i}\mathrm{d}\mu_{i}$, we arrive at the Gibbs–Duhem equation:
$v\mathrm{d}p+\sum_{i}\phi_{i}\mathrm{d}\mu_{i}=0,$ (113)
which has been often used to the formulate the pressure force for a unit volume
$\nabla p=\sum_{i}\phi_{i}\nabla\mu_{i}$ (114)
for the momentum evolution.
We note that a limitation of the above derivation is the omission of the kinetic energy, $\rho\mathbf{u}^{2}$ due to convection. The classical Gibbs–Duhem relation assumes thermodynamic equilibrium at each time step—that is, a quasi-static process—and therefore does not account for the momentum balance associated with kinetic energy. Another limitation arises in systems with large density ratios, where two variables, $\rho$ and $\phi$, are required to fully describe the change in system energy. In this context, the quintessential Gibbs–Duhem relation considers only the mechanical/chemical work associated with $\mathrm{d}\phi$, while the contribution from $\mathrm{d}\rho$ is neglected. A more complete treatment should consider the total energy, including internal energy ($\theta$), kinetic energy ($\rho\mathbf{u}\cdot\mathbf{u}$), and thermodynamic pressure ($p$), i.e., $\mathrm{d}(\theta+p+\rho\mathbf{u}\cdot\mathbf{u})$, rather than only $\mathrm{d}\theta$ as in the quasi-static approximation. When the contribution of the mixing entropy $s$ is considered, the internal energy $\theta$ should be replaced by the free energy density $f=\theta-Ts$. One illustrative example is that minimizing the sum of the thermodynamic pressure and kinetic energy naturally leads to momentum conservation:
$\mathrm{d}(p+\rho\mathbf{u}\cdot\mathbf{u})=\mathrm{d}p+\mathrm{d}(\rho\mathbf{u})\cdot\mathbf{u}+\rho\mathbf{u}\cdot\mathrm{d}\mathbf{u}.$ (115)
In this expression, the first two terms contribute to momentum conservation: $\mathrm{d}(\rho\mathbf{u})/\mathrm{d}t=-\nabla p$ by reformulating $\mathrm{d}p$ as $\mathrm{d}p/\mathrm{d}t=(\mathrm{d}p/\mathrm{d}\mathbf{r})\cdot(\mathrm{d}\mathbf{r}/\mathrm{d}t)=\nabla p\cdot\mathbf{u}$). The last term in Eq. (115) represents the dissipation of velocity. The comprehensive energy formulation encompassing the free energy, pressure energy, kinetic energy, and wall free energy is expressed in Eq. (52). The detailed derivation of the corresponding energy minimization, which represents a key result of the present work, is presented in Section VI.

### v3

@@ VI.2 Kinetic equations
The coefficient $\varsigma_{i}$ is the mobility. The physical meaning is that the diffusion leads to an increase in the mixing entropy $s_{\phi}$.
@@
The conservative force at the right hand side is responsible for the exchange of the potential energy $\mathcal{F}+\mathcal{P}$ with the kinetic energy $\mathcal{K}$. This result is nothing but momentum balance. Here, a more clear physical meaning of the force is presented; the force term consists of the pressure $p$ and the stress tensor $\underline{\underline{\bm{\Theta}}}$. The stress tensor is due to the gradient term in the potential energy $\mathcal{F}$, while the pressure is contributed by the associated free energy. Note that a viscous dissipation term cannot be added to the momentum conservation equation, as the pressure term does not contain the information of the velocity entropy [see Eq. (58)].
The only term left in the time evolution of system free energy, III, is $\int_{\Omega}\bm{\zeta}\cdot\dot{\mathbf{u}}\mathrm{d}\Omega$. By applying the same dissipation principle as done for the scalar field in Eq. (67), we obtain the dissipation equation for the velocity as
@@
Unlike the dissipation rule for the scalar variable $\phi_{i}$ with the associated potential $\nabla\mu_{i}$ [see Eq. (67)], three terms $\nabla\bm{\zeta}$, $\nabla\bm{\zeta}^{T}$, and $(\nabla\cdot\bm{\zeta})\mathbf{I}$ have to be taken into account for the dissipation of a vector $\mathbf{u}$. The physical meaning of the velocity dissipation is the increase of the velocity entropy $s_{u}$. Within the framework of entropy, the third term, III, in Eq. (65) can be interpreted as
$\bm{\zeta}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=-T\frac{\mathrm{d}s_{u}}{\mathrm{d}t}.$ (70)
The left hand side is responsible for the minimization of the kinetic energy, which is equivalent to the right hand side for the maximization of the velocity entropy.
@@
$\nabla\cdot\tau_{1}\nabla\bm{\zeta}=\nabla\cdot\tau_{1}\nabla(\partial k/\partial\mathbf{u});\ \bm{\zeta}=\partial k/\partial\mathbf{u};$ (72)
$\nabla\cdot\varsigma\nabla\mu=\nabla\cdot\varsigma\nabla(\partial e/\partial\mathbf{\phi});\ \mu=\partial e/\partial\phi.$ (73)
@@
The non-conserved form is known as the Allen-Cahn type dissipation following the gradient descent path, being consistent with the Langevin equation.
VI.3 Density evolution equation and the generalization of Bernoulli’s law
@@
As shown below, Eq. (75) provides a generalized form of Bernoulli’s law, a general definition of the sound speed, and a generalized equation of state (EOS).
@@
$\mathbf{u}(\mathbf{u}\cdot\nabla\rho)=-[\nabla\cdot(p\mathbf{I}+\underline{\underline{\bm{\Theta}}})].$ (76)
By considering a one-dimensional setup and using the relation $\mathrm{d}(\rho u^{2})=u^{2}\mathrm{d}\rho+2\rho u\mathrm{d}u$, Eq. (76) is further simplified as
$\mathrm{d}(p+\rho u^{2})=2\rho u\mathrm{d}u.$ (77)
This result is nothing but a generalized Bernoulli’s equation. When there is no change of the velocity entropy, namely, $\rho u\mathrm{d}u=0$, we have
@@
which is Bernoulli’s principle. However, when velocity entropy increases, $\mathrm{d}u\neq 0$, the associated energy is changed as well; in this case, the classic Bernoulli’s equation loses its validity.
VI.4 The problem of the conventional kinetic energy formulation
Conventional formulation for the kinetic energy is $\tilde{\mathcal{K}}=\int_{\Omega}\rho\mathbf{u}^{2}\mathrm{d}\Omega$. In this case, the time derivative of the kinetic energy reads
$\frac{\mathrm{d}{\tilde{\mathcal{K}}}}{\mathrm{d}t}=\int_{\Omega}\bigg(\frac{\mathrm{d}\rho}{\mathrm{d}t}\mathbf{u}^{2}+\rho\mathbf{u}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}\bigg)\mathrm{d}\Omega.$ (79)
A noteworthy shortcoming of Eq. (79) is the missing information of the momentum balance as well as the transformation between the potential and kinetic energies. Moreover, the first part in Eq. (79) requires a prior information about the time evolution of the density, which is unknown for two immiscible phases within the interface, as the excess volume within the interface is unknown. It is important to note that the classical continuity equation, $\partial_{t}\rho+\nabla\cdot(\rho\mathbf{u})=0$, cannot be applied here for density evolution. The reasons are three-fold:
(I) The classical continuity equation is derived for a one-component fluid, in which a single variable-either the density $\rho$ or the volume composition $\phi$-is sufficient to characterize the thermodynamic state. In such systems, density variations with pressure follow a standard equation of state (EOS), such as the van der Waals EOS. However, within the interface between two immiscible fluids, mixing entropy leads to a local excess volume that is generally non-zero and non-uniform in space. Under these conditions, the classical van der Waals EOS loses validity because it applies only to ideal mixing, where the excess volume of two mixing components is zero everywhere. For non-ideal mixing of two components, three volume contributions must be considered: the component volumes $v_{1}$ and $v_{2}$, and the excess volume $v_{e}$. When $v_{e}=0$, the classical EOS may be applied to $v_{1}$ and $v_{2}$, relating the partial pressures $p_{1}$ and $p_{2}$ to the corresponding densities $\rho_{1}$ and $\rho_{2}$. When $v_{e}\neq 0$, an additional EOS governing $v_{e}$ is required, and this cannot be handled within the classical framework of continuity equations because an additional degree of freedom due to excess volume has been introduced. As discussed in Section III, our approach is to redistribute the excess volume $v_{e}$ into $v_{1}$ and $v_{2}$ and to introduce a new volume concentration that incorporates the excess-volume information. With this redistribution, the effective densities $\rho_{1}$ and $\rho_{2}$ are no longer constants; their actual values, including the redistributed excess volume, depend on the chosen redistribution scheme and are not of primary interest here.
Instead, we derive a new density equation that directly relates density variations to changes in both pressure and velocity entropy [Eq. (75)]. This stands in contrast to the classical continuity equation, in which density variations are independent of pressure and entropy. In the classical framework, one must prescribe an EOS externally to obtain the kinetic evolution. In the present theory, by contrast, the EOS emerges naturally from the evolution equations themselves (see Section VIB), aided by the pressure obtained from the modified Gibbs–Duhem relation and the newly defined volume concentration that incorporates the excess volume.
(II) In the case of a spatially uniform density for one-component fluids, the classical continuity equation reduces to $\nabla\cdot\mathbf{u}=0$. This divergence-free condition, which characterizes incompressible one-component fluids, does not hold within the interface between two immiscible fluids. Enforcing $\nabla\cdot\mathbf{u}=0$ in such regions produces a pressure field that lacks physical meaning. When the excess volume is nonzero, the fluid is inherently compressible, and thus deviates from the incompressibility assumption. In this case, the pressure associated with the compressibility arising from the excess volume must be determined from the modified Gibbs–Duhem relation, rather than from solving $\nabla\cdot\mathbf{u}=0$. As already discussed in Section III, two variables are required to describe the thermodynamic state of the fluids. The volume concentration accounts for the compressibility and determines the pressure according to Eq. (58). The density evolution equation then naturally gives rise to EOS.
(III) The classical continuity equation cannot incorporate diffusion processes, such as evaporation and condensation, which are incompatible with the incompressibility assumption, as noted in the Introduction and discussed elsewhere Abels et al. (2012). In fact, the diffusion equation arises from the presence of excess volume or equivalently from the mixing entropy $s_{\phi}$, thereby leading to an effective compressibility.
@@ VI.5 Replication of the Navier-Stokes equations
$\frac{1}{2}\rho\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=\rho\{\nabla\cdot\tau_{1}[\nabla(\rho\mathbf{u})+\nabla(\rho\mathbf{u})^{T}]+\nabla\cdot\tau_{2}\nabla\cdot(\rho\mathbf{u})\mathbf{I}\}=\nabla\cdot\eta(\nabla\mathbf{u}+\nabla\mathbf{u}^{T})+\nabla\cdot[\nu(\nabla\cdot\mathbf{u})\mathbf{I}].$ (82)
@@
We stress that viscous dissipation inherently contradicts the incompressibility condition. When viscous dissipation occurs, density cannot remain constant, as dissipative processes introduce entropy changes that modify the density. However, under dissipation-free conditions (e.g., isentropic flow), density can reasonably be treated as constant. In this way, the Navier-Stokes equation has the conflicts between viscous dissipation and invariance of density (incompressibility). In fact, Eq. (81) cannot be summed with Eq. (82) because:
@@ VI.6 Remarks
The key point of the present theory is the energy minimization based on an alternative formulation for the kinetic energy, $\rho\mathbf{u}\cdot\mathbf{u}$, rather than the conventional formulation, $\rho\mathbf{u}^{2}$. In our derivation, the dissipation of the kinetic energy is divided into two parts. The first part is the momentum conservation, $\mathrm{d}(\rho\mathbf{u})/\mathrm{d}t$ which changes not only the fluid velocity $\mathbf{u}$ but also the density $\rho$. This derivation differs from Euler and Navier-Stokes equations where the momentum variation is only caused by the evolution of the fluid velocity. The second part of the kinetic energy dissipation is the velocity dissipation associated with the velocity entropy.
In the current work, we consider the linear momentum $\bm{\zeta}=\rho\mathbf{u}$. Non-linear terms related to fluid velocity may be accounted for the dissipation if we consider non-linear momentum, for example, $\bm{\zeta}\propto\mathbf{u}^{n}$ or $\bm{\zeta}\propto(\nabla\mathbf{u})^{n}$. The non-linear momentum not only modifies the velocity dissipation but also changes the momentum balance, which may be applied for understanding highly nonlinear effects. These specialized cases are out of the scope of the present discussion.
@@
The evolution of the velocity corresponds to the dissipation according to Eq. (69). The local velocity can increase (acceleration) or decrease (deceleration), but overall the kinetic energy $\mathcal{K}$ plus other potential energies decreases with time in a closed system. Once dissipation takes place, i.e., $\mathrm{d}\mathbf{u}/\mathrm{d}t\neq 0$, the density at the interface cannot be assumed to be a constant; in this case, the density evolves with time. The density evolution is especially important when the fluid velocity is close to the speed of sound, i.e., high Mach number fluid. Note that in reality, dissipation cannot be avoided, since entropy always increases with time.
@@
The first part corresponds to the momentum balance. The second part is associated with the velocity dissipation.
@@ VII.1 Sound speed
which is consistent with the results of the perturbation analysis Landau and Lifshitz (2013). Note that two conditions have been applied to derive the sound speed: $\delta q\approx 0$ and $\eta\approx 0$, known as isentropic condition (adiabatic and reversible). For one component fluid, the Landau potential is expressed as $f-\sum_{i}\mu_{i}\phi_{i}=f$, where the free energy $f$ is written as $f=\theta-Ts$. In the case of isentropic condition, the free energy is mainly contributed by the internal energy, which consists of microscopic kinetic energy $\varrho$ and microscopic potential energy $\xi$,
@@ VII.2 Equation of the state (EOS)
which is the EOS of ideal gas with a factor, $\ln\rho$. Note that the thermodynamic pressure $\Psi$ should be used instead of $p$. The term $\ln\rho$ is analogous to the entropy contribution, and this result is consistent with previous work on the free energy formulation of a water-gas system Ledesma-Aguilar et al. (2014). By dividing both sides of Eq. (97) by the molar volume, we obtain $\Psi\sim\rho\ln\rho$, which is the prescribed EOS used in Refs. Ledesma-Aguilar et al. (2014); Ridl and Wagner (2018). This highlights our central result: the density evolution in our theory naturally generates the EOS, in contrast to the prior EOS imposed in existing diffuse-interface methods.
@@ VII.3 Water-air interface
$u^{2}\frac{\mathrm{d}\rho}{\mathrm{d}x}=-\frac{\mathrm{d}p}{\mathrm{d}x},$ (106)
@@ Acknowledgements.
This research is supported by VirtMat project of the Helmholtz association, as part of the programme “MSE-materials science and engineering” number-43.31.01. F.W. conceived the overall concept, conducted all theoretical derivations, and prepared the manuscript. B.N. contributed to securing the funding.
Appendix A The modified Gibbs-Duhem equation and the pressure
The modified Gibbs-Duhem relation reads
$\delta p+\sum_{i=1}^{K}\phi_{i}\delta\mu_{i}+\rho\mathbf{u}\cdot\delta\mathbf{u}=0.$ (113)
The velocity term is replaced by the entropy contribution as $\rho\mathbf{u}\cdot\delta\mathbf{u}=-T\delta s_{u}=-\delta(Ts_{u})$. In the last equality, we have used the isothermal condition. The chemical potential term $\phi_{i}\delta\mu_{i}$ is rewritten as $\phi_{i}\delta\mu_{i}=\delta(\mu_{i}\phi_{i})-\mu_{i}\delta\phi_{i}$ according to the chain rule of total derivative. Recalling the definition of the chemical potential, $\mu_{i}=\delta e/\delta\phi_{i}$, we further write the chemical potential term $\phi_{i}\delta\mu_{i}$ as $\sum_{i=1}^{K}\phi_{i}\delta\mu_{i}=\sum_{i=1}^{K}\delta(\mu_{i}\phi_{i})-\delta e$ (see Appendix C). Summarizing all these equalities, we obtain the following equation
$\delta(-p+\sum_{i=1}^{K}\mu_{i}\phi_{i}-e-Ts_{u})=\delta(-p+\sum_{i=1}^{K}\mu_{i}\phi_{i}-e_{0})=0.$ (114)
In the above equation, we have used the equality $e+Ts_{u}=e_{0}$. Integrating Eq. (114), we obtain the final expression for the pressure as
$p=e_{0}-\sum_{i=1}^{K}\mu_{i}\phi_{i}+p_{\text{ref}},$ (115)
where $p_{\text{ref}}$ is a constant. For simplicity, this constant is set as zero in the current work.
Appendix B $\delta$ operator, $\nabla$ operator, and the formulation of the chemical potential
If replacing the $\delta$ operator by the space gradient operator $\nabla$ and the velocity term by the velocity entropy $T\nabla s_{u}$ in Eq. (113), the following equation for the gradient of the pressure is obtained
$\nabla p=-\sum_{i=1}^{K}\phi_{i}\nabla\mu_{i}+T\nabla s_{u}.$ (116)
When using the original definition of the chemical potential as
$\mu_{i}=\frac{\mathrm{d}e}{\mathrm{d}\phi_{i}}=\partial e/\partial\phi_{i}+(\partial e/\partial\nabla\phi_{i})\cdot\nabla\phi_{i},$ (117)
which is consistent with Ref. Ridl and Wagner (2018), Eq. (116) is rewritten as
$\nabla p-\nabla(Ts_{u})+\sum_{i=1}^{K}[\nabla(\mu_{i}\phi_{i})-\mu_{i}\nabla\phi_{i}]=0.$ (118)
Integrating Eq. (118) by using Eq. (117), we obtain the same formulation of Eq. (115).
Note that if the conventional functional derivative
$\mu_{i}=\partial e/\partial\phi_{i}-\kappa_{i}\nabla^{2}\phi_{i}=\partial f_{0}/\partial\phi_{i}-\kappa_{i}\nabla^{2}\phi_{i}$ (119)
is used Ledesma-Aguilar et al. (2014), one cannot obtain Eq. (115). By using Eq. (119), one obtains a different formulation of the pressure $p=f_{0}-\sum_{i=1}^{K}\kappa_{i}(\nabla\phi_{i})^{2}-\sum_{i=1}^{K}\mu_{i}\phi_{i}+p_{\text{ref}}$; a negative sign occurs in the second term. The reason for this inconsistency is as follows. Only after applying the integration by parts to the term $(\partial e/\partial\nabla\phi_{i})\cdot\nabla\phi_{i}$ in Eq. (117) on the whole domain $\Omega$, we obtain the second term with a negative sign in Eq. (119). This negative sign arises from integration by parts and leads to inconsistent results for the pressure. One cannot use this result—obtained after applying integration by parts in the whole domain—to integrate the Gibbs–Duhem relation, which is a local differential equation. Instead, the original definition given in Eq. (117) should be used when integrating the Gibbs–Duhem relation.
Appendix C Derivative of $\mu\phi$
Supposing that $\phi(x)$ and $\phi(y)$ are functions: $\mathbb{R}^{3}\mapsto\mathbb{R}$, and $\mu$ is a functional: $\Phi\mapsto\mathbb{R}$, where $\phi(x),\ \phi(y)\in\Phi$, the derivative for the product of $\mu$ and $\phi$ reads Gelfand et al. (2000)
$\frac{\delta[\mu\phi(x)]}{\delta\phi(y)}=\phi(x)\frac{\delta\mu}{\delta\phi(y)}+\mu\frac{\delta\phi(x)}{\delta\phi(y)}.$ (120)
In a special case, $\phi(x)=\phi(y)$, we have $\frac{\delta\phi(x)}{\delta\phi(y)}=1$, and therefore $\delta(\mu\phi)=\phi\delta\mu+\mu\delta\phi$. As $\mu=\delta e/\delta\phi$, the above equation is finally expressed as $\delta(\mu\phi)=\phi\delta\mu+\delta e$.